A 2.5 kg banana falls from a bunch that is 3.5 m above the ground. Calculate the speed with which the banana will strike the ground.


If the banana makes a tiny 0.050 m imprint in the mud, what force must the ground exert on the banana?


1
Expert's answer
2021-02-15T17:44:32-0500

1) According to the conservation of energy law, the initial potential energy of the banana


"W= mgh"

is equal to its final kinetic energy


"K = \dfrac{mv^2}{2}"

Here "m = 2.5kg" is the mass, "g = 9.8m\/s" is the gravitational acceleration, "h = 3.5m" is the initial height, "v" is the speed with which the banana will strike the ground. Expressing "v", obtain:


"mgh = \dfrac{mv^2}{2}\\
v = \sqrt{2gh}\\
v = \sqrt{2\cdot 3.5\cdot 9.8} \approx 8.28 m/s"

2) According to the work-energy theorem, the change in the kinetic energy of the banana is equal to the work experienced by it from the mud. The kinetic energy has changed from


"K_1 = \dfrac{mv^2}{2}"

to "K_2" . In turn, the work is defined as follows:


"A = F_{net}s"

where "F_{net }" is the force the ground exerted on the banana, and "s = 0.05m" is the distance travelled by the banana in the mud. Thus, obtain:


"\dfrac{mv^2}{2} = F_{net}s\\
F_{net} = \dfrac{mv^2}{2s} = \dfrac{2mgh}{2s} = \dfrac{mgh}{s}\\
F_{net} =\dfrac{2.5\cdot 9.8\cdot 3.5}{0.05} = 1715N"

Answer. 1) 8.28 m/s, 2) 1715 N.